40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 244     Entry time: Thu Jan 17 14:13:20 2008
Author: rob 
Type: Update 
Category: LSC 
Subject: Wednesday's locking 
Incremental progress on locking yet again. This time the handoff of DARM to the OMC worked, and progress halted at handing off control of the common mode to REFL166.
ELOG V3.1.3-